OpenStudy (jamesreed30):

Find the exact area of a circle having the given circumference 8 PI

OpenStudy (blazeryder):

\[C=2 \times \pi \times r\]\[8\pi=2 \times r\]\[8\pi \div 2 = 2r \div2\]\[r=4\]

OpenStudy (blazeryder):

Now what would we do next?

OpenStudy (jamesreed30):

4 times 4 =16

OpenStudy (blazeryder):

Then?

OpenStudy (jamesreed30):

16 times 3.14

OpenStudy (blazeryder):

Yup! Now what is the answer?

OpenStudy (jamesreed30):

Which I get 50.24 but my computer says this is wrong

OpenStudy (blazeryder):

Let me see something real quick

OpenStudy (blazeryder):

Try 16pi

OpenStudy (blazeryder):

It said to find the exact so keep it in terms of pi.

OpenStudy (blazeryder):

@Jamesreed30

OpenStudy (mathmale):

Here you have to work backwards, since you are given the circumference of a circle and need to find the radius. Once you have the radius, you can find the area of the circle. What is the formula for the circumference of a circle of radius r? Solve it for the variable r.